OpenStudy (anonymous):

How would y – x2 = 5 be written in function notation?

OpenStudy (haseeb96):

y=5+x^2

OpenStudy (anonymous):

f(x)=x2+5

OpenStudy (anonymous):

or as Haseeb said if you meant x squared

OpenStudy (haseeb96):

y=f(x) so f(x)=x^2 +5